# Approximate Sequential Optimization for Informative Path Planning

### v1

@@ 6.3 A and D-IPP Results
    MCTS: the belief MDP approach presented by [17] [17].
@@
Examples of the trajectories produced by each of the methods are shown in Figure 4. In these example trajectories, we have also included the multimodal sensing aspect of the problem introduced in Section 5.3. MCTS used the method described by [17] to select locations for the high-quality sensors [17]. The random method selected high-quality sensor locations from its action space and was limited to placing $k=3$ high-quality sensors. The other methods all used the convex formulation provided in Section 5.3. From the trajectories, we can see that the random solution is the worst performer and stays near the starting location (bottom left corner) until it reaches the budget limit and must head to the goal location leaving much of the environment unexplored. The greedy method slightly improves upon the random method, but is unable to coordinate its actions several steps into the future therefore leading to inefficient exploration. MCTS improves upon the greedy solution; however, much of its exploration is limited to small areas around where it has already visited. The exact and $\textbf{tr}(\Sigma^{-1})$ methods are able to better distribute the measurement locations resulting in more efficient exploration. Lastly, the approximate sequential path optimization method tries to fully cover the environment by distributing its measurements accordingly and outperforms the exact and $\textbf{tr}(\Sigma^{-1})$ methods in terms of the best (lowest) value of the A-IPP objective.
@@
Figure 5 compares the multimodal sensor selection methods of [17] and the method we present in Section 5.3. For these experiments, we started with the paths that were produced from the MCTS policy. To construct these paths, the MCTS policy chooses from an action space that includes both movement and sensing actions. Therefore, each point along the trajectory includes a location and sensor type to be used. To compare the sensor selection methods, we left the MCTS path solution constant and used this as the initial path for the convex multimodal sensor selection problem. We then evaluated both sensor selection methods on the A and D-IPP objectives (i.e. the paths were fixed for both methods, but the sensor types were changed). Figure 5 shows that the convex multimodal sensor selection method we propose outperforms that of MCTS. Note that these comparisons were done with $k=3$ high-quality sensors.
@@ 6.6 Local Optimization
As discussed by [1] and [24], we can perform a local optimization step to greedily improve our approximate solution [1, 24]. This local optimization procedure is the polish step in the relax-round-polish method presented by [24] [24]. Starting from the approximate solution $\hat{z}$ presented in Section 4, we check other paths that can be derived from $\hat{z}$ by swapping one of the $k$ nodes along the path with one of the one-hop neighbors available. In this case, we define a valid one-hop neighbor of node $i$ to be a node $j$ such that $j\in N_{out}(i)$ and if node $i$ were removed from the path and node $j$ were added then the total length of the path would not change. Equivalently, if $l_{A}=\sum_{i=1}^{n}\sum_{j\in N_{out}(i)}\hat{z}_{ij}$ represents the total length of path $A$ before a valid swap occurs then $l_{A}$ must equal $l_{B}$ where path $B$ contains a valid swap derived from path $A$.
@@ 6.7 Other Objectives
Figure 10 provides an example of the ASPO method applied to a multi-agent IPP scenario with $M=3$ agents in addition to obstacles being present in the environment. Each agent $A_{i}$ is tasked with traversing the graph to collect measurements. The agents share their measurements, forming a collective measurement vector. We can see that the agents are able to distribute their exploration throughout the environment to effectively split up the environment into thirds leading to more efficient exploration. The solution to this multi-agent problem using our ASPO method was computed in under $5$ seconds on a graph with $n=400$ nodes.
@@ 7 Conclusion
Future work includes incorporating uncertainty in the localization of the agent as it traverses the graph. Introducing and managing this uncertainty can better mirror real-world situations where precise localization is challenging. Another future extension to consider would be adding additional constraints such as the refueling variant of the problem where there are various fuel stations located in the environment and the agent must also plan where to fill up at each station. Extending this to problems where the fuel costs are uncertain would also result in emergent behaviors that are relevant to many real-world domains. We also wish to further extend the multi-agent version of the problem to consider communication constraints, intra-agent action inference, and unknown obstacles. These extensions would increase the combinatorial nature of the problem, likely necessitating new approximate methods to allow for similar scalability.

### v2

@@ 6.3 A and D-IPP Results
    MCTS: the belief MDP approach presented by Ott et al. [22].
@@
Examples of the trajectories produced by each of the methods are shown in figure 4. In these example trajectories, we have also included the multimodal sensing aspect of the problem introduced in section 5.3. MCTS used the method described by Ott et al. to select locations for the high-quality sensors [22]. The random method selected high-quality sensor locations from its action space and was limited to placing $k=3$ high-quality sensors. The other methods all used the convex formulation provided in section 5.3. From the trajectories, we can see that the random solution is the worst performer and stays near the starting location (bottom left corner) until it reaches the budget limit and must head to the goal location leaving much of the environment unexplored. The greedy method slightly improves upon the random method, but is unable to coordinate its actions several steps into the future therefore leading to inefficient exploration. MCTS improves upon the greedy solution; however, much of its exploration is limited to small areas around where it has already visited. The exact and $\textbf{tr}(\Sigma^{-1})$ methods are able to better distribute the measurement locations resulting in more efficient exploration. Lastly, the approximate sequential path optimization method tries to fully cover the environment by distributing its measurements accordingly and outperforms the exact and $\textbf{tr}(\Sigma^{-1})$ methods in terms of the best (lowest) value of the A-IPP objective.
@@
Figure 5 compares the multimodal sensor selection methods of Ott et al. and the method we present in section 5.3. For these experiments, we started with the paths that were produced from the MCTS policy. To construct these paths, the MCTS policy chooses from an action space that includes both movement and sensing actions. Therefore, each point along the trajectory includes a location and sensor type to be used. To compare the sensor selection methods, we left the MCTS path solution constant and used this as the initial path for the convex multimodal sensor selection problem. We then evaluated both sensor selection methods on the A and D-IPP objectives (i.e. the paths were fixed for both methods, but the sensor types were changed). Figure 5 shows that the convex multimodal sensor selection method we propose outperforms that of MCTS. Note that these comparisons were done with $k=3$ high-quality sensors.
@@ 6.6 Local Optimization
As discussed by Joshi and Boyd and Diamond et al., we can perform a local optimization step to greedily improve our approximate solution [1, 29]. This local optimization procedure is the polish step in the relax-round-polish method presented by Diamond et al. [29]. Starting from the approximate solution $\hat{z}$ presented in section 4, we check other paths that can be derived from $\hat{z}$ by swapping one of the $k$ nodes along the path with one of the one-hop neighbors available. In this case, we define a valid one-hop neighbor of node $i$ to be a node $j$ such that $j\in N_{out}(i)$ and if node $i$ were removed from the path and node $j$ were added then the total length of the path would not change. Equivalently, if $l_{A}=\sum_{i=1}^{n}\sum_{j\in N_{out}(i)}\hat{z}_{ij}$ represents the total length of path $A$ before a valid swap occurs then $l_{A}$ must equal $l_{B}$ where path $B$ contains a valid swap derived from path $A$.
@@ 6.7 Other Objectives
Figure 10: Examples of trajectories planned by the CAtNIPP trajectory sampling variant and the ASPO method on a graph with $400$ nodes and available budget of $6$.
Method Budget 4 Budget 6 Budget 8 Budget 10 Budget 12
$\mathbf{tr}(\Sigma)$ T(s) $\mathbf{tr}(\Sigma)$ T(s) $\mathbf{tr}(\Sigma)$ T(s) $\mathbf{tr}(\Sigma)$ T(s) $\mathbf{tr}(\Sigma)$ T(s)
CG 19.67($\pm$1.57) 0.15 5.42($\pm$0.29) 0.19 2.57($\pm$0.14) 0.26 1.58($\pm$0.09) 0.33 1.24($\pm$0.07) 0.42
CTS 15.94($\pm$1.18) 44.04 4.94($\pm$0.23) 62.35 2.27($\pm$0.12) 88.78 1.54($\pm$0.08) 111.88 1.24($\pm$0.07) 143.94
ASPO 14.02($\pm$0.87) 0.69 4.79($\pm$0.28) 1.26 2.24($\pm$0.13) 1.96 1.28($\pm$0.11) 2.61 0.98($\pm$0.11) 3.33
Table 1: Comparison between CAtNIPP Greedy (CG), CAtNIPP Trajectory Sampling (CTS), and ASPO IPP solvers on the AIPP problem using the environment setup presented by Cao et al. with $400$ nodes, $20$ edge connections, and $4$ trajectory samples for CTS [30]. The average covariance matrix trace $\mathbf{tr}(\Sigma)$ in high interest areas is reported after running out of budget (standard error in parentheses). T(s) is the average total planning time in seconds. Results from $100$ simulation runs for each budget.
6.8 Comparison with Attention-based Network Approach
We also compared our approximate method with the Context-Aware Attention-based Network for Informative Path Planning (CAtNIPP) approach presented by Cao et al. [30]. CAtNIPP is a state-of-the-art deep reinforcement learning based approach using attention-based neural networks to learn policies that balance short-term exploitation and long-term exploration for the AIPP problem.
For even comparison between our two methods, we used the graph and environment setup described by Cao et al. [30]. Specifically, a Matérn $3/2$ Kernel with length scale of $0.45$. The ground truth environment is constructed by averaging $8$ to $12$ random 2-dimensional Gaussian distributions in the unit square $[0,1]^{2}$, to construct the true interest map [30].
The objective is to reduce the variance in high-interest areas defined as regions that have true quantities of interest greater than or equal to $0.4$. To effectively solve this variant of the AIPP problem, solution methods must balance the tradeoff between collecting additional measurements in high-interest areas they have previously discovered with exploring new areas of the environment they have not yet visited. Cao et al. present two different variants of CAtNIPP. The greedy variant works in the standard reinforcement learning manner where the agent selects the action with the highest activation in its policy. The trajectory sampling variants plan a $15$-step trajectory and execute the first $3$ steps before replanning.
Table 1 presents the results across several different budgets. We see that across all of the budgets, our ASPO method consistently achieves the lowest variance in high-interest areas showing that it more effectively balances the exploration exploitation tradeoff. We see that the CAtNIPP Greedy variant has the lowest planning time across all of the methods considered but also is consistently the worst performer. ASPO is quite close to the Greedy variant in terms of planning time, while the trajectory sampling variant requires significantly more planning time. Figure 10 also shows examples of trajectories planned by the CAtNIPP trajectory sampling variant and ASPO in addition to the posterior estimates after executing the trajectories.
It is important to note that this comparison was performed on the specific environment that CAtNIPP was trained on. CAtNIPP is sensitive to changes in the hyperparameters of the Gaussian process kernel function and the upper bound confidence that defines the high-interest areas. Therefore, CAtNIPP would require retraining if any of these parameters were changed significantly. Cao et al. report that it took around $24$ hours training on four GPUs to converge. Therefore, one of the key strengths of ASPO over neural network based approaches is that no retraining is required when any of the Gaussian process hyperparameters, graph structures, budgets, objectives, or confidence intervals are changed.
6.9 Multi-agent Example
Figure 11 illustrates the application of the ASPO method in a multi-agent Informative Path Planning (IPP) scenario, where $M=3$ agents are deployed in an environment that includes obstacles. Each agent $A_{i}$ is assigned the task of traversing through the environment to collect measurements. The measurements gathered by the individual agents are then shared among them, resulting in a collective measurement vector that represents the combined data from all agents.
We can see that the agents are able to distribute their exploration throughout the environment to effectively split up the environment into thirds leading to more efficient exploration. Notably, the solution to this multi-agent problem, computed using our ASPO method, was achieved in less than $5$ seconds, on a graph with $400$ nodes. This highlights the efficiency and scalability of the ASPO method, particularly in complex environments with multiple agents and obstacles.
@@ 7 Conclusion
7.1 Limitations and Future Work
While our method has proven effective, there are some key limitations that present opportunities for future research. First, our approach assumes that the location of the agent is fully observable, the current energy state of the agent is known, and the amount of budget used by specific actions is deterministic. Incorporating uncertainty in these aspects would make our method more robust and better aligned with real-world scenarios where localization errors, energy consumption variability, and unexpected environmental changes are common. Additionally, our current work assumes full knowledge of these variables, and introducing uncertainty in these areas could lead to more sophisticated and resilient planning strategies.
Another promising area for future research involves exploring more complex multi-agent scenarios. Specifically, we aim to investigate cases where communication between agents is limited or unreliable, which is a common challenge in real-world applications. Additionally, introducing adversarial agents into the problem space, which may attempt to disrupt or compete with the planning agents, could add another layer of complexity and realism to the problem. Addressing these challenges will likely require the development of new approximate methods that can maintain scalability while handling the increased combinatorial complexity introduced by these factors.
@@ Acknowledgments
Appendix A Additional Experiments
We also compared our method against a solution method that first solves the convex relaxation presented in section 3.1. Notice that the solution to the convex relaxation does not produce a valid path; however, we can construct a valid path by rounding the solution. The rounding procedure starts at the starting node and then selects the neighboring node with the minimum $u$ value that is greater than the current nodes $u$ value given by the constraints in equation 11. The results on the A and D-IPP problems are shown in figure 12 and figure 13.
Figure 12: Runtime and A-IPP objective as a function of graph size for ASPO and the relax and rounded solution. Each curve shows the average runtime and objective value respectively with the standard error reported over 25 different simulations. The dashed line on the left indicates the 120 second runtime constraint imposed on all methods.
Figure 13: Runtime and D-IPP objective as a function of graph size for ASPO and the relax and rounded solution. Each curve shows the average runtime and objective value respectively with the standard error reported over 25 different simulations. The dashed line on the left indicates the 120 second runtime constraint imposed on all methods.
Appendix B Derivation of Maximum a Posteriori Estimate
  • Here we derive the maximum a posteriori estimate of $x$ given in equation equation 3. Given the linear measurement model $y_{k}=a_{i_{k}}^{T}x+\nu_{i_{k}}$, where $a_{i}$ is a known vector, $x$ is the parameter vector we want to estimate, and $\nu_{i}$ is additive Gaussian noise with zero mean and variance $\sigma^{2}$, the likelihood of observing a particular $y_{i}$ collected along a path with node indices $(i_{1},i_{2},\dots,i_{L})$ given $x$ is:
    $p(y_{i}|x)\propto\exp\left(-\frac{(y_{i}-a_{i_{k}}^{T}x)^{2}}{2\sigma^{2}}\right).$ (35)
    The likelihood of observing the measurements $y=[y_{1},y_{2},\dots,y_{L}]^{T}$ given $x$ is:
    $p(y|x)=\prod_{k=1}^{L}p(y_{k}|x)=\prod_{k=1}^{L}\mathcal{N}(y_{k}|a_{i_{k}}^{T}x,\sigma^{2})$ (36)
    Since the noise is Gaussian, this can be written as:
    $p(y|x)=\left(\frac{1}{\sqrt{2\pi\sigma^{2}}}\right)^{L}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{k=1}^{L}(y_{k}-a_{i_{k}}^{T}x)^{2}\right)$ (37)
    The prior distribution over $x$ is assumed to be Gaussian, $\mathcal{N}(\bar{x},\Sigma_{x})$:
    $p(x)\propto\exp\left(-\frac{1}{2}(x-\bar{x})^{T}\Sigma_{x}^{-1}(x-\bar{x})\right).$ (38)
    From Bayes’ rule, the posterior distribution $p(x|y)$ is:
    $p(x|y)\propto p(y|x)p(x)$ (39)
    Substituting the expressions for the likelihood and the prior:
    $p(x|y)\propto\exp\left(-\frac{1}{2\sigma^{2}}\sum_{k=1}^{L}(y_{k}-a_{i_{k}}^{T}x)^{2}\right)\cdot\exp\left(-\frac{1}{2}(x-\bar{x})^{T}\Sigma_{x}^{-1}(x-\bar{x})\right)$ (40)
The MAP estimate $\hat{x}$ maximizes the posterior probability. Equivalently, it minimizes the negative log-posterior probability:
$\hat{x}=\operatorname*{argmin}_{x}\left[\frac{1}{2\sigma^{2}}\sum_{k=1}^{L}(y_{k}-a_{i_{k}}^{T}x)^{2}+\frac{1}{2}(x-\bar{x})^{T}\Sigma_{x}^{-1}(x-\bar{x})\right]$ (41)
The solution can be found by setting the gradient with respect to $x$ to zero:
$\frac{\partial}{\partial x}\left[\frac{1}{2\sigma^{2}}\sum_{k=1}^{L}(y_{k}-a_{i_{k}}^{T}x)^{2}+\frac{1}{2}(x-\bar{x})^{T}\Sigma_{x}^{-1}(x-\bar{x})\right]=0$ (42)
$-\frac{1}{\sigma^{2}}\sum_{k=1}^{L}a_{i_{k}}(y_{k}-a_{i_{k}}^{T}x)+\Sigma_{x}^{-1}(x-\bar{x})=0$ (43)
Simplifying:
$\frac{1}{\sigma^{2}}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}x-\frac{1}{\sigma^{2}}\sum_{k=1}^{L}a_{i_{k}}y_{k}+\Sigma_{x}^{-1}x-\Sigma_{x}^{-1}\bar{x}=0$ (44)
$\left(\frac{1}{\sigma^{2}}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)x=\frac{1}{\sigma^{2}}\sum_{k=1}^{L}a_{i_{k}}y_{k}+\Sigma_{x}^{-1}\bar{x}$ (45)
Rearranging:
$\hat{x}=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}y_{k}+\Sigma_{x}^{-1}\bar{x}\right)$ (46)
Distributing the first term:
$\hat{x}=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}y_{k}\right)+\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\Sigma_{x}^{-1}\bar{x}$ (47)
We can rewrite the second term as follows:
$\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\Sigma_{x}^{-1}\bar{x}$ (48)
$=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}\bar{x}-\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}\bar{x}+\Sigma_{x}^{-1}\bar{x}\right)$
$=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\left(-\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}\bar{x}+\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)\bar{x}\right)$
$=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\left(-\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}\bar{x}\right)+\bar{x}$
Substituting this back into equation 47:
$\hat{x}$ $=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}y_{k}\right)$ (49)
$+\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\left(-\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}\bar{x}\right)+\bar{x}$
$=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\left(\sigma^{-2}\sum_{k=1}^{L}\left(a_{i_{k}}y_{k}-a_{i_{k}}a_{i_{k}}^{T}\bar{x}\right)\right)+\bar{x}$
Recall $\bar{y}_{k}=a_{i_{k}}^{T}\bar{x}$:
$\hat{x}=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}y_{k}-a_{i_{k}}\bar{y}_{k}\right)+\bar{x}$ (50)
Finally:
$\hat{x}=\bar{x}+\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\sigma^{-2}\sum_{k=1}^{L}(y_{k}-\bar{y}_{k})a_{i_{k}}$ (51)
Appendix C Derivation of Posterior Covariance
  • The covariance of the posterior distribution is the inverse of the Hessian of the negative log-posterior with respect to $x$, evaluated at the MAP estimate. The Hessian can be found starting from the expression for the negative log-posterior
    $-\log p(x|y)=\frac{1}{2\sigma^{2}}\sum_{k=1}^{L}(y_{k}-a_{i_{k}}^{T}x)^{2}+\frac{1}{2}(x-\bar{x})^{T}\Sigma_{x}^{-1}(x-\bar{x})+\text{const}$ (52)
    Computing the Hessian:
    $H=\frac{\partial^{2}}{\partial x^{2}}\left(\frac{1}{2\sigma^{2}}\sum_{k=1}^{L}(y_{k}-a_{i_{k}}^{T}x)^{2}+\frac{1}{2}(x-\bar{x})^{T}\Sigma_{x}^{-1}(x-\bar{x})\right)$ (53)
    The second derivative of the first term is:
    $\frac{\partial^{2}}{\partial x^{2}}\left(\frac{1}{2\sigma^{2}}\sum_{k=1}^{L}(y_{k}-a_{i_{k}}^{T}x)^{2}\right)=\frac{1}{\sigma^{2}}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}$ (54)
    The second derivative of the second term is:
    $\frac{\partial^{2}}{\partial x^{2}}\left(\frac{1}{2}(x-\bar{x})^{T}\Sigma_{x}^{-1}(x-\bar{x})\right)=\Sigma_{x}^{-1}$ (55)
    Thus, the Hessian is:
    $H=\frac{1}{\sigma^{2}}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}$ (56)
    The covariance matrix of the posterior distribution is the inverse of the Hessian:
    $\Sigma=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}$ (57)
which is equation 4. ∎
Appendix D Gaussian Process Equivalence
A Gaussian process is a distribution over functions that can be used to predict an underlying function $f$ given some previously observed noisy measurements at location x. That is, $y=f(\textbf{x})+z$ where $f$ is deterministic but $z$ is zero-mean Gaussian noise, $z\sim\mathcal{N}(0,\sigma^{2})$. The distribution of functions conditioned on the previously observed measurements $\bm{\hat{y}}\mid\bm{y},\sigma^{2}\sim\mathcal{N}(\bm{\mu^{\ast}},\bm{\Sigma^{\ast}})$ is given by:
$\bm{\mu^{\ast}}=\bm{m}(X^{\ast})\newline +\bm{K}(X^{\ast},X)(\bm{K}(X,X)+\sigma^{2}\bm{I})^{-1}(\bm{y}-\bm{m}(X))$ (58)
$\bm{\Sigma^{\ast}}=\bm{K}(X^{\ast},X^{\ast})-\bm{K}(X^{\ast},X)(\bm{K}(X,X)+\sigma^{2}\bm{I})^{-1}\bm{K}(X,X^{\ast})$ (59)
where $X$ is the set of previously measured locations, $X^{\ast}$ is the set of locations we wish to predict the values $\bm{\hat{y}}$, $\bm{m}(X)$ is the mean function and $\bm{K}(X,X^{\prime})$ is the covariance matrix constructed with the kernel $k(\bm{x},\bm{x^{\prime}})$ [24].
We want to show that equation 58 and equation 59 are equivalent to
$\hat{x}=\bar{x}+\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}\sigma^{-2}\sum_{k=1}^{L}(y_{k}-\bar{y}_{k})a_{i_{k}},$ (60)
and
$\Sigma=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}$ (61)
from equation 3 and equation 4 respectively.
D.1 Covariance Equivalence
  • We know that each $a_{i}\in\mathbf{R}^{m}$ characterizes the measurements and that there are $n$ unique $a_{i}$ corresponding to each of the $n$ locations in the environment. In the Gaussian process formulation, $X$ denotes the previously measured locations and therefore corresponds to the path indices $p=(i_{1},i_{2},\dots,i_{L})$. Then
    $A=\begin{bmatrix}-a_{i_{1}}^{T}-\\ -a_{i_{2}}^{T}-\\ \vdots\\ -a_{i_{L}}^{T}-\end{bmatrix}$ (62)
    with $A\in\mathbf{R}^{L\times m}$. Notice that $\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}=A^{T}A$ with $A^{T}A\in\mathbf{R}^{m\times m}$.
For equivalence, we have that the query points $X^{\ast}$ correspond to the locations in the environment that we want to estimate the vector $x$ at. [^3] Notice that $X^{\ast}$ correspond to locations in the environment while $x$ represents the vector we wish to estimate, i.e. the values at those locations in the environment. Then we have the following equalities:
  1. 1.
    $m(X^{\ast})=\bar{x}$
  2. 2.
    $\bm{K}(X^{\ast},X^{\ast})=\Sigma_{x}$
  3. 3.
    $\bm{K}(X^{\ast},X)=\Sigma_{x}A^{T}$
where the last equality is specified by the sensor characterization.
Using the Woodbury matrix identity [^4] The Woodbury matrix identity: $\left(A+UCV\right)^{-1}=A^{-1}-A^{-1}U\left(C^{-1}+VA^{-1}U\right)^{-1}VA^{-1}$ we can rewrite equation 59 as
$\bm{\Sigma^{\ast}}=\left(\bm{K}(X^{\ast},X^{\ast})^{-1}+\bm{K}(X^{\ast},X^{\ast})^{-1}\bm{K}(X^{\ast},X)\sigma^{-2}\bm{I}\bm{K}(X,X^{\ast})\bm{K}(X^{\ast},X^{\ast})^{-1}\right)^{-1}$ (63)
Substituting in equalities $2$ and $3$:
$\bm{\Sigma^{\ast}}$ $=\left(\Sigma_{x}^{-1}+\sigma^{-2}\Sigma_{x}^{-1}\bm{K}(X^{\ast},X)\bm{K}(X,X^{\ast})\Sigma_{x}^{-1}\right)^{-1}$ (64)
$=\left(\Sigma_{x}^{-1}+A^{T}A\right)^{-1}$
$=\left(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1}\right)^{-1}$
$=\Sigma$
D.2 Mean Equivalence
  • Starting from the expression for the Gaussian process mean and using the previous result from Woodbury matrix identity [^5] Notice we had $\bm{K}(X,X)=VA^{-1}U=\bm{K}(X,X^{\ast})\bm{K}(X^{\ast},X^{\ast})^{-1}\bm{K}(X^{\ast},X)$ in the Woodbury matrix identity.:
    $\bm{\mu^{\ast}}$ $=\bm{m}(X^{\ast})+\bm{K}(X^{\ast},X)(\bm{K}(X,X)+\sigma^{2}\bm{I})^{-1}(\bm{y}-\bm{m}(X))$ (65)
    $=\bm{m}(X^{\ast})\newline +\bm{K}(X^{\ast},X)(\bm{K}(X,X^{\ast})\bm{K}(X^{\ast},X^{\ast})^{-1}\bm{K}(X^{\ast},X)+\sigma^{2}\bm{I})^{-1}(\bm{y}-\bm{m}(X))$
    Substituting in $\bm{K}(X^{\ast},X)=\Sigma_{x}A^{T}$ and $\bm{K}(X,X^{\ast})=A\Sigma_{x}$:
    $\bm{\mu^{\ast}}$ $=\bm{m}(X^{\ast})+\Sigma_{x}A^{T}(A\Sigma_{x}\Sigma_{x}^{-1}\Sigma_{x}A^{T}+\sigma^{2}\bm{I})^{-1}(\bm{y}-\bm{m}(X))$ (66)
    $=\bm{m}(X^{\ast})+\Sigma_{x}A^{T}(A\Sigma_{x}A^{T}+\sigma^{2}\bm{I})^{-1}(\bm{y}-\bm{m}(X))$
    Notice that $\Sigma_{x}A^{T}(A\Sigma_{x}A^{T}+\sigma^{2}\bm{I})^{-1}=(A^{T}\sigma^{-2}\bm{I}A+\Sigma_{x}^{-1})^{-1}A^{T}\sigma^{-2}\bm{I}$. To see this, multiply
    $\Sigma_{x}A^{T}(A\Sigma_{x}A^{T}+\sigma^{2}\bm{I})^{-1}\stackrel{{\scriptstyle?}}{{=}}(A^{T}\sigma^{-2}\bm{I}A+\Sigma_{x}^{-1})^{-1}A^{T}\sigma^{-2}\bm{I}$ (67)
    on the left by $(A^{T}\sigma^{-2}\bm{I}A+\Sigma_{x}^{-1})$ and on the right by $(A\Sigma_{x}A^{T}+\sigma^{2}\bm{I})$ to get
    $(A^{T}\sigma^{-2}\bm{I}A+\Sigma_{x}^{-1})\Sigma_{x}A^{T}\stackrel{{\scriptstyle?}}{{=}}A^{T}\sigma^{-2}\bm{I}(A\Sigma_{x}A^{T}+\sigma^{2}\bm{I})\$ (68)
    which is true.
    Now substituting in $(A^{T}\sigma^{-2}\bm{I}A+\Sigma_{x}^{-1})^{-1}A^{T}\sigma^{-2}\bm{I}$ for $\Sigma_{x}A^{T}(A\Sigma_{x}A^{T}+\sigma^{2}\bm{I})^{-1}$ in equation 66 we get:
    $\bm{\mu^{\ast}}$ $=\bm{m}(X^{\ast})+(A^{T}\sigma^{-2}\bm{I}A+\Sigma_{x}^{-1})^{-1}A^{T}\sigma^{-2}\bm{I}(\bm{y}-\bm{m}(X))$ (69)
    $=\bar{x}+(\sigma^{-2}A^{T}A+\Sigma_{x}^{-1})^{-1}\sigma^{-2}A^{T}(\bm{y}-\bm{m}(X))$
    $=\bar{x}+(\sigma^{-2}\sum_{k=1}^{L}a_{i_{k}}a_{i_{k}}^{T}+\Sigma_{x}^{-1})^{-1}\sigma^{-2}\sum_{k=1}^{L}(y_{k}-\bar{y}_{k})a_{i_{k}}$
    $=\hat{x}$
Appendix E Mutual Information
Starting from the Gaussian process formulation provided by equation 58 and equation 59 we have that the distribution of functions conditioned on the previously observed measurements $\bm{\hat{y}}\mid\bm{y},\sigma^{2}\sim\mathcal{N}(\bm{\mu^{\ast}},\bm{\Sigma^{\ast}})$ is given by:
$\bm{\mu^{\ast}}=\bm{m}(X^{\ast})\newline +\bm{K}(X^{\ast},X)(\bm{K}(X,X)+\sigma^{2}\bm{I})^{-1}(\bm{y}-\bm{m}(X))$ (70)
$\bm{\Sigma^{\ast}}=\bm{K}(X^{\ast},X^{\ast})-\bm{K}(X^{\ast},X)(\bm{K}(X,X)+\sigma^{2}\bm{I})^{-1}\bm{K}(X,X^{\ast}).$ (71)
For a multivariate normal distribution $\mathbf{z}\sim\mathcal{N}(0,\mathbf{P})$ where $z\in\mathbf{R}^{n}$, the differential entropy is given by:
$H(\mathbf{z})=\frac{1}{2}\log\left((2\pi e)^{n}\det(\mathbf{P})\right).$ (72)
For the conditional distribution $\hat{\mathbf{y}}\mid\mathbf{y}$, which is Gaussian with covariance $\bm{\Sigma}^{*}$, the conditional entropy is:
$H(\hat{\mathbf{y}}\mid\mathbf{y})=\frac{1}{2}\log\left((2\pi e)^{m}\det(\bm{\Sigma}^{*})\right)$ (73)
since we are interested in making predictions at the $m$ locations in the environment [31].
The mutual information between the observed outputs $\mathbf{y}$ and the predicted outputs $\hat{\mathbf{y}}$ is given by:
$I(\hat{\mathbf{y}};\mathbf{y})=H(\hat{\mathbf{y}})-H(\hat{\mathbf{y}}\mid\mathbf{y}).$ (74)
Given that $\hat{\mathbf{y}}$ and $\mathbf{y}$ have a joint Gaussian distribution, the marginal entropy $H(\hat{\mathbf{y}})$ can be computed similarly:
$H(\hat{\mathbf{y}})=\frac{1}{2}\log\left((2\pi e)^{m}\det(\mathbf{K}(X^{*},X^{*}))\right)$ (75)
Therefore, the mutual information is:
$I(\hat{\mathbf{y}};\mathbf{y})=\frac{1}{2}\log\left(\frac{\det(\mathbf{K}(X^{*},X^{*}))}{\det(\bm{\Sigma}^{*})}\right)=\frac{1}{2}\log\left(\det(\mathbf{K}(X^{*},X^{*}))\right)-\frac{1}{2}\log\left(\det(\bm{\Sigma}^{*})\right).$ (76)
Rewriting this using the notation from section 2 we have:
$I(\hat{x};\{y_{1},\dots,y_{L}\})=\frac{1}{2}\log\left(\frac{\det(\Sigma_{x})}{\det(\Sigma)}\right)=\frac{1}{2}\log\left(\det(\Sigma_{x})\right)-\frac{1}{2}\log\left(\det(\Sigma)\right).$ (77)
Therefore maximizing the mutual information is equivalent to minimizing $\log\left(\det(\Sigma)\right)$.